OpenStudy (anonymous):

Solve x² + 7x + 2 = 0

OpenStudy (anonymous):

you mean x² + 7x + 12 = 0 ????????

OpenStudy (jackellyn):

Use the quadratic equation: a=1 b=7 c=2 \[-7 \pm \sqrt{7^{2}-4(1)(2)}/(2(1))=-7 \pm \sqrt{49-8}/2=-7 \pm \sqrt{41}/2\]

OpenStudy (anonymous):

Here, a = 1, b = 7 and c = 12 As: \[D = b^2 - 4ac\] So, from here can you find D???

OpenStudy (jackellyn):

how do you know it is a 12?

OpenStudy (anonymous):

Sorry, Little mistake.. c = 2

OpenStudy (anonymous):

i just thought is it 12 ???

OpenStudy (anonymous):

yes or no ???

OpenStudy (anonymous):

no